OpenStudy (pphalke):

A fraternity is to elect a president and a treasurer from the group of 40 members. How many ways can these two officers be elected?

OpenStudy (yanasidlinskiy):

The number of ways to elect a president: \[C \frac{ 1 }{ 40 } = 40\] \[C= \frac{ 1 }{ 39 } = 39\] \[Possible Combinations: N=40*39=1560\] So, two officers can be elected with1560 different ways.

OpenStudy (nincompoop):

what is C?